Skip to main contentSkip to solution

Probabilities to solve a specific problem by A, B and C are 12,13\frac{1}{2}, \frac{1}{3} and 14\frac{1}{4} respectively. Probability that at least one will solve the problem is:

Solution

✅ Correct Option: 4

P(none solves)=12⋅23⋅34=14P(\text{none solves}) = \frac{1}{2} \cdot \frac{2}{3} \cdot \frac{3}{4} = \frac{1}{4}. So P(at least one solves)=1−14=34P(\text{at least one solves}) = 1 - \frac{1}{4} = \frac{3}{4}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question